OpenStudy (anonymous):

In an arithmetic sequence, a5=-1+9i and a11=7+21i. Find the value of a30.

sam (.sam.):

Do you know what's the formula for AP?

OpenStudy (anonymous):

No, I dont know the formula.

sam (.sam.):

The formula is \[a_n=a_1+(n-1)d\] n=term, d=common difference, so, we have \(a_5\) and \(a_{11}\) \[a_5=-1+9i \\ \\ a_1+(5-1)d=-1+9i.....(1) \\ \\ --------------- \\ \\ a_{11}=7+21i \\ \\ a_1+(11-1)d=7+21i.....(2)\] Solve for \(a_1\) and \(d\) here

OpenStudy (anonymous):

Thank you!